# Complete regularity of linear cocycles and the Baire category of the set of Lyapunov-Perron regular points

### v1

@@ 4.5. All-or-meager results
Let $g:Y\to Y$ be a homeomorphism of a compact space admitting a fully-supported Borel probability measure. Let $\mathcal{B}$ be a linear cocycle over $g$, with generator $B$, and let $\mathcal{R}_{\mathcal{B}}$ be the set of LP-regular points. We can easily arrange examples so that $\mathcal{R}_{\mathcal{B}}\neq Y$ (using Theorem 4.8, for instance). On the other hand, there exists a point $z\in\mathcal{R}_{\mathcal{B}}$ whose alpha- and omega-limits are both equal to $Y$. That is, if $z_{n}=g^{n}(z)$, then for any $N>0$, the sets $\{z_{n}:n>N\}$ and $\{z_{n}:n<-N\}$ are both dense in $Y$. Now fix an increasing two-sided sequence $(t_{n})_{n\in\mathbb{Z}}$ such that $t_{n}\to 0$ as $n\to-\infty$ and $t_{n}\to 1$ as $n\to+\infty$. Consider the following space:
$X\coloneqq(Y\times\{0,1\})\cup\{(z_{n},t_{n}):n\in\mathbb{Z}\},$
which is a compact subset of $Y\times[0,1]$. Define a homeomorphism $f\colon X\to X$ as follows:
$f(y,t)\coloneqq(g(y),t)\text{ if }t=0,1,\quad f(z_{n},t_{n})\coloneqq(z_{n+1},t_{n+1})\,.$
Finally, define a cocycle $\mathcal{A}$ over $f$ whose generator is $\mathcal{A}(y,t)\coloneqq B(y)$. Then the set $\mathcal{R}_{\mathcal{A}}$ of LP-regular points is not the whole space $X$, but it is neither a meager set, because it contains the orbit of $(z_{0},t_{0})$, which is an open and dense set.
@@
Letting $f$ be a transitive Anosov diffeomorphism again, it is an open problem whether Theorem 4.8 remains true if we assume that the cocycle is only continuous, instead of being Hölder continuous.
@@
Theorem 4.10 extends some previous results such as [25, Theorem 4] and [30, Corollary 1.4].
Johnson and Zampogni asked a question in [33, p. 104], which in our terminology is as follows: Assume that the dynamics $f$ is minimal, and that for all points $x$ and all nonzero vectors $v$, the forward Lyapunov exponent $\lim_{n\to\infty}\frac{1}{n}\log\|\mathcal{A}(x,n)v\|$ is well-defined; does it follow that the cocycle is completely regular? Note that Theorem 4.10 provides a positive answer to this question under the stronger hypothesis that all points are LP-regular.
@@ 4.6. Singular values of cocycles
Each singular value $\sigma_{k}(L)$ depends continuously on $L$. The extreme singular values are the norm $\|L\|=\sigma_{1}(L)$ and the co-norm $m(L)=\sigma_{d}(L)$. More generally, by the Courant–Fischer principle [29, Theorem 7.3.8], for any $k=1,\dots,d$,
[…]
@@
The list of singular values of the exterior power $L^{\wedge r}$ coincides with the list $\sigma_{k_{1}}(L)+\cdots+\sigma_{k_{j}}(L)$, where $1\leq k_{1}<\cdots<k_{r}\leq d$, with repetitions being taken into account.
@@
$\mathcal{A}(f^{j}(x),n)=\mathcal{A}(f^{n}(x),j)\,\mathcal{A}(x,n)\,\mathcal{A}(T^{j}x,-j)\,.$
@@
Let $f$ be a uniquely ergodic homeomorphism of a compact metric space and let $\mathcal{A}$ be a continuous $\mathrm{SL}^{\pm}(2,\mathbb{R})$-cocycle over $f$. If the cocycle $\mathcal{A}$ is not uniformly hyperbolic, then there exists a residual subset $G_{0}$ of $X$ such that for all $x\in G_{0}$, the set of accumulation points of the sequence $\frac{1}{n}\log\|\mathcal{A}(x,n)\|$ is the interval $[0,\chi_{1}]$, where $\chi_{1}=\chi_{1}(\mu)$ is the first Lyapunov exponent of the cocycle with respect to the unique $f$-invariant measure $\mu$.
Note that the result above improves some of the conclusions of [25, Theorem 4].
@@
For each $x$, let $\Lambda_{x}$ be the set of accumulation points of the sequence $\frac{1}{n}\log\|\mathcal{A}(x,n)\|$. Then $\Lambda_{x}$ is a compact interval (see e.g. [58, Lemma 2.5]). Note that $\Lambda_{x}\subset[0,\infty)$ because $|\det\mathcal{A}|\equiv 1$, and $\Lambda_{x}\subset[0,\chi_{1}]$ due to Proposition 2.3. If $\chi_{1}=0$, there is nothing to show. So, assume that $\chi_{1}$ is positive. Since the cocycle is not uniformly hyperbolic, by Theorem 4.12 there is a dense set of points $x$ for which lim $\liminf\frac{1}{n}\log\|\mathcal{A}(x,n)\|=0$. On the other hand, there is another dense set of points $x$ for which $\lim\frac{1}{n}\log\|\mathcal{A}(x,n)\|=\chi_{1}$. The proof of Theorem 4.1 actually shows that if $0<\alpha^{\prime\prime}<\beta^{\prime\prime}<\chi_{1}$, then there is a residual subset of points $x$ such that
lim $\liminf\frac{1}{n}\log\|\mathcal{A}(x,n)\|<\alpha^{\prime\prime}<\beta^{\prime\prime}<\limsup\frac{1}{n}\log\|\mathcal{A}(x,n)\|$
@@
Let $f$ be a minimal homeomorphism $f$ of a compact metric space. Let $\mathcal{A}$ be a $d$-dimensional continuous linear cocycle over $f$ and let $\mathcal{R}$ be the set of its LP-regular points. Suppose that $\mathcal{A}$ admits no dominated splitting into non-trivial subbundles. Then:
@@
whose determinants are ±$\pm 1$. Our assumption ensures that all Lyapunov exponents of $\hat{\mathcal{A}}$ with respect to all ergodic measures are zero. By Proposition 3.2, the cocycle $\hat{\mathcal{A}}$ is completely regular with zero spectrum.
@@
By minimality of $f$, the measure $\mu$ has full support, so the formula above holds for a dense set of points $y$. On the other hand, by Theorem 4.12, there is another dense set where the formula fails, and actually the liminf of the sequence is 00. Now we can apply Theorem 4.1 and conclude that the set of points $z$ for which
@@ 5.1. All points being regular does not imply complete regularity
If we take a continuous cocycle of diagonal matrices, then every point will be LP-regular. But the Lyapunov exponents will not be constant, in general. This example was pointed out before in [33, Example 4.1].
Let us exhibit an example where $f$ is topologically transitive. Consider the group $G\coloneqq\mathrm{SL}(2,\mathbb{R})$, the lattice $\Gamma\coloneqq\mathrm{SL}(2,\mathbb{Z})$, and the quotient $N\coloneqq{{}_{\Gamma}}\backslash^{G}$, that is, the set of right cosets $\Gamma g$. Then $N$ is a Hausdorff but non-compact space with respect to the quotient topology (actually it is the unit tangent bundle of the modular surface). Let $X\coloneqq N\cup\{\infty\}$ be the one-point compactification of $N$. Define $f\colon X\to X$ by $f(\infty)\coloneqq\infty$ and
@@
The map $f$ is a homeomorphism and, like the previous example (5.1), is such the Birkhoff averages of every continuous function converge pointwise (see [20, Theorem 5.1]); furthermore, $f$ is topologically transitive. Arguing as before, Condition (1) does not imply Condition (2) for this map $f$.
@@
    A co-compact lattice $\Gamma\subset G$ such that $\Gamma\cap\mathrm{SL}(2,\mathbb{R})$ is a co-compact lattice in $\mathrm{SL}(2,\mathbb{R})$ (see [59, Exerc. 4, §18.6]).
@@
    The quotient $N\coloneqq{{}_{\Gamma}}\backslash^{G}$, that is, the set of right cosets $\Gamma g$; this is a compact smooth manifold of dimension $8$; we denote by $m_{N}$ the normalized volume in $N$ which is induced on $N$ by the Haar measure on the group $G$.
@@
By Ratner’s equidistribution theorem [48, Theorem B], for every $x\in N$, the orbit $\{\varphi^{n}x\}_{n\geq 0}$ is equidistributed with respect to some probability measure $m_{x}$ on $N$, that is, for every continuous function $h\colon N\to\mathbb{R}$, we have
[…]
@@
Let $\widetilde{G}$ be the subgroup of $G$ consisting of the matrices of the form $\left(\begin{smallmatrix}*&*&0\\ *&*&0\\ 0&0&1\end{smallmatrix}\right)$, that is, the standard copy of $\mathrm{SL}(2,\mathbb{R})$ in $\mathrm{SL}(3,\mathbb{R})$. Let $\pi\colon G\to N$ be the quotient map $\pi(g)\coloneqq\Gamma g$ and let $N_{0}\coloneqq\pi(\widetilde{G})$. Since the unipotent subgroup $(u^{t})$ is contained in $\widetilde{G}$, the set $N_{0}$ is invariant under the unipotent flow $\{\varphi^{t}\}$. Recall that our lattice $\Gamma<G$ was chosen so that $\widetilde{\Gamma}\coloneqq\Gamma\cap\widetilde{G}$ is a co-compact lattice in $\widetilde{G}$. Therefore, the quotient $\widetilde{N}\coloneqq{{}_{\widetilde{\Gamma}}}\backslash^{\widetilde{G}}$ is a compact $3$-manifold which is actually the unit tangent bundle of a compact hyperbolic surface. We define a unipotent flow on $\widetilde{N}$ by the formula $\widetilde{\varphi}^{t}(\widetilde{\Gamma}\widetilde{g})\coloneqq\widetilde{\Gamma}\widetilde{g}u^{t}$; this is actually the horocycle flow on the unit tangent bundle. Define a map $\widetilde{N}\to N_{0}$ by $\widetilde{\Gamma}\widetilde{g}\mapsto\Gamma\widetilde{g}$. This map is a homeomorphism, and it conjugates the flows $\widetilde{\varphi}^{t}$ and $\varphi^{t}|_{N_{0}}$. By a theorem of Furstenberg, the horocycle flow $\widetilde{\varphi}^{t}$ is strictly ergodic (i.e., minimal and uniquely ergodic); furthermore, the time-$1$ map $\widetilde{\varphi}^{1}$ is also strictly ergodic: see [22]. We conclude that the submanifold $N_{0}$ of $N$ supports a unique $\varphi^{1}$-invariant measure, which we call $m_{N_{0}}$.
@@
Note that $a^{s}\not\in\widetilde{G}$ for all nonzero $s$. Since the projection $\pi\colon G\to N$ is a covering map and $\pi(\widetilde{G})$ is the compact set $N_{0}$, it follows that $\pi(a^{s})\not\in N_{0}$ for all sufficiently small $s\neq 0$. But $\pi(a_{s})\in N_{s}$, and so $N_{s}\neq N_{0}$ for all sufficiently small $s\neq 0$. Since $N_{s}$ is the support of the measure $m_{N_{s}}$, this proves that the curve $s\mapsto m_{N_{s}}$ is non-constant, as we wanted to show. ∎
@@
Note that the element $(\alpha,1)$ of the $\mathbb{R}\times\mathbb{Z}$-action is the diffeomorphism $f_{\alpha}$. Let $A$ be the set of $\alpha\in\mathbb{R}_{+}$ such that $(f_{\alpha},m_{N}\times m_{S})$ is not ergodic. By a result of Pugh and Shub [46, Theorem 2], the set $A$ is countable. Since the measure $m_{N}\times m_{S}$ is positive on open sets, the map $f_{\alpha}$ is topologically transitive whenever it is ergodic. ∎
@@
Equality (5.6) follows from the definition of the map $f_{\alpha}$, and the following two facts: (a) all the Lyapunov exponents of $\varphi^{t}$ are zero; and (b) the Lyapunov exponent of the Anosov flow in its flow direction is zero. Equality (5.9) follows from (5.4) and the fact that the function $\tau$ is bounded. To prove equality (5.7) observe that for every $n>0$,
@@ 5.2. Agreement of Lyapunov spectra does not imply complete regularity
If $f\colon X\to X$ is a homeomorphism of a compact metric space $X$, we say that $f$ is a Veech map if it is strictly ergodic (i.e. minimal and uniquely ergodic) and $f^{2}$ is minimal but not uniquely ergodic. Veech maps were first constructed by Veech [57].
Walters [58] has shown that given a Veech map $f$, one can construct a non-uniformly hyperbolic cocycle with base $f$. We will construct a specific Walters cocycle with some additional properties which will be instrumental in our proof of Theorem 5.5. For this cocycle we need the base Veech map constructed in [9, Sec. 4] (the idea of this construction is due to Scott Schmieding).
@@
[…]
where $\overline{e_{k}}$ denotes the word obtained from $e_{k}$ by switching the arrows. Let $X$ be the subset of $\mathsf{A}^{\mathbb{Z}}$ formed by those bi-infinite sequences $\omega=(\omega_{i})_{i\in\mathbb{Z}}$ such that for every $n>0$, the word $\omega_{-n}\omega_{-n+1}\cdots\omega_{n}$ occurs as a subword (i.e., a factor) of some $e_{k}$. According to [9, Theorem 4.1], $X$ is a closed, shift-invariant subset of $\mathsf{A}^{\mathbb{Z}}$, and if $f\colon X\to X$ is the restricted shift, then $f$ is a Veech map. Define a continuous matrix valued map
@@
If $\omega$ is a LP-regular point, then the two sequences above must converge to the same limit as $n\to\infty$. However, according to [9, Lemma 4.12], this common limit, whenever it exists, is a number $c>0$ independent of $\omega$. Therefore all LP-regular points of our Walters cocycle have the same Lyapunov exponents $c$ and $-c$.

### v2

@@ 4.5. All-or-meager results
Let $g:Y\to Y$ be a homeomorphism of a compact metric space $Y$ and let $\mathcal{B}$ be a linear cocycle over $g$, with generator $B$. We assume two conditions: First, $g$ admits a fully-supported Borel probability measure $\mu$. Second, the set $\mathcal{R}_{\mathcal{B}}$ of LP-regular points is not the whole space $Y$. Examples satisfying these conditions can be found using Theorem 4.8, for instance. Note that $\mu$-almost every point in $Y$ has alpha- and omega-limit sets both equal to $Y$. Fix a point $y_{0}$ with these properties which is also an element of the full-measure set $\mathcal{R}_{\mathcal{B}}$. Let $y_{n}\coloneqq g^{n}(y_{0})$; then, for any $N>0$, the sets $\{y_{n}:n>N\}$ and $\{y_{n}:n<-N\}$ are both dense in $Y$. Now fix an increasing two-sided sequence $(t_{n})_{n\in\mathbb{Z}}$ such that $t_{n}\to 0$ as $n\to$-$\infty$ and $t_{n}\to 1$ as $n\to+\infty$. Define the following subsets of $Y\times[0,1]$:
$Z\coloneqq\{(y_{n},t_{n}):n\in\mathbb{Z}\}\quad\text{and}\quad X\coloneqq(Y\times\{0,1\})\cup Z\,.$
Then $\overline{Z}=X$ and in particular $X$ is compact. Also note that each point in $Z$ has a neighborhood disjoint from $Y\times\{0,1\}$ and in particular $Z$ is a relatively open subset of $X$. Define a map $f\colon X\to X$ as follows:
$f(y,t)\coloneqq(g(y),t)\text{ if }t=0,1,\quad f(y_{n},t_{n})\coloneqq(y_{n+1},t_{n+1})\,.$
Then $f$ is actually a homeomorphism of $X$, and the subset $Z$ is the orbit of $(y_{0},t_{0})$. Finally, define a cocycle $\mathcal{A}$ over $f$ whose generator is $\mathcal{A}(y,t)\coloneqq B(y)$. Then the set $\mathcal{R}_{\mathcal{A}}$ of LP-regular points is not the whole space $X$, as its complement includes $(Y\setminus\mathcal{R}_{\mathcal{B}})\times\{0,1\}$. On the other hand, since the set $\mathcal{R}_{\mathcal{B}}$ contains the point $y_{0}$, the set $\mathcal{R}_{\mathcal{A}}$ contains the point $(y_{0},t_{0})$, as well as its orbit $Z$. Since $Z$ is an open and dense subset of $X$, it follows that $\mathcal{R}_{\mathcal{A}}$ is not a meager set. Thus, we have exhibited an example where the all-or-meager dichotomy fails.
@@
Returning to the case of transitive Anosov diffeomorphisms, it is an open problem whether Theorem 4.8 extends to cocycles that are merely continuous rather than Hölder continuous.
@@
Theorem 4.10 extends some previous results such as [27, Theorem 4] and [32, Corollary 6.6].
Johnson and Zampogni asked a question in [35, p. 104], which in our terminology is as follows: Assume that the dynamics $f$ is minimal, and that for all points $x$ and all nonzero vectors $v$, the forward Lyapunov exponent $\lim_{n\to\infty}\frac{1}{n}\log\|\mathcal{A}(x,n)v\|$ is well-defined; does it follow that the cocycle is completely regular? Note that Theorem 4.10 provides a positive answer to this question under the stronger hypothesis that all points are LP-regular.
@@ 4.6. Singular values of cocycles
Each singular value $\sigma_{k}(L)$ depends continuously on $L$. The extreme singular values are the norm $\|L\|=\sigma_{1}(L)$ and the co-norm $m(L)=\sigma_{d}(L)$. More generally, by the Courant–Fischer principle [31, Theorem 7.3.8], for any $k=1,\dots,d$,
$\sigma_{k}(L)$ $=\max\big{\{}m(A|_{E})$ $\colon$ $\text{$E$ is a subspace of dim.\ $k$}\big{\}}$
$=\min\big{\{}\|A|_{F}\|$ $\colon$ $\text{$F$ is a subspace of dim.\ $d-k+1$}\big{\}}\,.$
@@
The list of singular values of the exterior power $L^{\wedge r}$ coincides with the list $\sigma_{k_{1}}(L)+\cdots+\sigma_{k_{r}}(L)$, where $1\leq k_{1}<\cdots<k_{r}\leq d$, with repetitions being taken into account.
@@
$\mathcal{A}(f^{j}(x),n)=\mathcal{A}(f^{n}(x),j)\,\mathcal{A}(x,n)\,\mathcal{A}(f^{j}(x),-j)\,.$
@@
Let $f$ be a uniquely ergodic homeomorphism of a compact metric space and let $\mathcal{A}$ be a continuous $\mathrm{SL}^{\pm}(2,\mathbb{R})$-cocycle over $f$. If the cocycle $\mathcal{A}$ is not uniformly hyperbolic, then there exists a residual subset $G_{0}$ of $X$ such that for all $x\in G_{0}$, the set of accumulation points of the sequence $\frac{1}{n}\log\|\mathcal{A}(x,n)\|$ is the interval $0,\chi_{1}]$, where $\chi_{1}=\chi_{1}(\mu)$ is the first Lyapunov exponent of the cocycle with respect to the unique $f$-invariant measure $\mu$.
Note that the result above improves some of the conclusions of [27, Theorem 4].
@@
For each $x$, let $\Lambda_{x}$ be the set of accumulation points of the sequence $\frac{1}{n}\log\|\mathcal{A}(x,n)\|$. Then $\Lambda_{x}$ is a compact interval (see e.g. [63, Lemma 2.5]). Note that $\Lambda_{x}\subset[0,\infty)$ because $|\det\mathcal{A}|\equiv 1$, and $\Lambda_{x}\subset[0,\chi_{1}]$ due to Proposition 2.3. If $\chi_{1}=0$, there is nothing to show. So, assume that $\chi_{1}$ is positive. Since the cocycle is not uniformly hyperbolic, by Theorem 4.12 there is a dense set of points $x$ for which $\liminf\frac{1}{n}\log\|\mathcal{A}(x,n)\|=0$. On the other hand, there is another dense set of points $x$ for which $\lim\frac{1}{n}\log\|\mathcal{A}(x,n)\|=\chi_{1}$. The proof of Theorem 4.1 actually shows that if $0<\alpha^{\prime\prime}<\beta^{\prime\prime}<\chi_{1}$, then there is a residual subset of points $x$ such that
$\liminf\frac{1}{n}\log\|\mathcal{A}(x,n)\|<\alpha^{\prime\prime}<\beta^{\prime\prime}<\limsup\frac{1}{n}\log\|\mathcal{A}(x,n)\|$
@@
Let $f$ be a minimal homeomorphism of a compact metric space $X$. Let $\mathcal{A}$ be a $d$-dimensional continuous linear cocycle over $f$ and let $\mathcal{R}$ be the set of its LP-regular points. Suppose that $\mathcal{A}$ admits no dominated splitting into non-trivial subbundles. Then:
@@
whose determinants are $\pm 1$. Our assumption ensures that all Lyapunov exponents of $\hat{\mathcal{A}}$ with respect to all ergodic measures are zero. By Proposition 3.2, the cocycle $\hat{\mathcal{A}}$ is completely regular with zero spectrum.
@@
By minimality of $f$, the measure $\mu$ has full support, so the formula above holds for a dense set of points $y$. On the other hand, by Theorem 4.12, there is another dense set where the formula fails, and actually the liminf of the sequence is 0. Now we can apply Theorem 4.1 and conclude that the set of points $z$ for which
@@ 5.1. All points being regular does not imply complete regularity
If we take a continuous cocycle of diagonal matrices, then every point will be LP-regular. But the Lyapunov exponents will not be constant, in general. This example was pointed out before in [35, Example 4.1].
[…]
@@
The map $f$ is a homeomorphism and, like the previous example (5.1), is such the Birkhoff averages of every continuous function converge pointwise (see [21, Theorem 5.1]); furthermore, $f$ is topologically transitive. Arguing as before, Condition (1) does not imply Condition (2) for this map $f$.
@@
    A co-compact lattice $\Gamma\subset G$ such that $\Gamma\cap\mathrm{SL}(2,\mathbb{R})$ is a co-compact lattice in $\mathrm{SL}(2,\mathbb{R})$ (see [65, Exerc. 4, §18.6]).
@@
[…]
@@
By Ratner’s equidistribution theorem [52, Theorem B], for every $x\in N$, the future and past orbits $\{\varphi^{\pm n}x\}_{n\geq 0}$ are equidistributed with respect to a probability measure $m_{x}$ on $N$, that is, for every continuous function $h\colon N\to\mathbb{R}$, we have
(5.5) $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}h(\varphi^{\pm j}x)=\int_{N}h\,dm_{x}\,.$
Furthermore (see [52, p. 255]), the measures $m_{x}$ are ergodic, and thus $\varphi$ is pointwise ergodic in the terminology of [23].
@@
[…]
@@
We have constructed a continuous curve $s\mapsto m_{N_{s}}$ in the space of $\varphi^{1}$-invariant Borel probability measures on $N$ (endowed with the weak topology). We will prove that this curve is non-constant. Once we do that, the lemma will immediately follow, as it will be sufficient to take any $C^{\infty}$ strictly positive function $\tau$ for which the function $s\mapsto\int_{N}\tau\,dm_{N_{s}}$ is non-constant.
@@
Note that the element $(\alpha,1)$ of the $\mathbb{R}\times\mathbb{Z}$-action is the diffeomorphism $f_{\alpha}$. Let $A$ be the set of $\alpha\in\mathbb{R}_{+}$ such that $(f_{\alpha},m_{N}\times m_{S})$ is not ergodic. By a result of Pugh and Shub [50, Theorem 2], the set $A$ is countable. Since the measure $m_{N}\times m_{S}$ is positive on open sets, the map $f_{\alpha}$ is topologically transitive whenever it is ergodic. ∎
@@
Equality (5.6) follows from the definition of the map $f_{\alpha}$, and the following two facts: (a) all the Lyapunov exponents of $\varphi^{t}$ are zero; and (b) the Lyapunov exponent of the Anosov flow in its flow direction is zero. Equality (5.9) follows from (5.4) and the fact that the function $\tau$ is bounded. To prove equality (5.7), observe that for every $n>0$,
@@ 5.2. Agreement of Lyapunov spectra does not imply complete regularity
If $f\colon X\to X$ is a homeomorphism of a compact metric space $X$, we say that $f$ is a Veech-like map if it is strictly ergodic (i.e. minimal and uniquely ergodic) and $f^{2}$ is minimal but not uniquely ergodic. Veech-like maps were first constructed by Veech [62].
Walters [63] has shown that given a Veech-like map $f$, one can construct a non-uniformly hyperbolic cocycle with base $f$. We will construct a specific Walters cocycle with some additional properties which will be instrumental in our proof of Theorem 5.5. For this cocycle we need the base Veech-like map constructed in [9, Sec. 4] (the idea of this construction is due to Scott Schmieding).
@@
where $\overline{e_{k}}$ denotes the word obtained from $e_{k}$ by switching the arrows. Let $X$ be the subset of $\mathsf{A}^{\mathbb{Z}}$ formed by those bi-infinite sequences $\omega=(\omega_{i})_{i\in\mathbb{Z}}$ such that for every $n>0$, the word $\omega_{-n}\omega_{-n+1}\cdots\omega_{n}$ occurs as a subword (i.e., a factor) of some $e_{k}$. According to [9, Theorem 4.1], $X$ is a closed, shift-invariant subset of $\mathsf{A}^{\mathbb{Z}}$, and if $f\colon X\to X$ is the restricted shift, then $f$ is a Veech-like map. Define a continuous matrix valued map
@@
If $\omega$ is a LP-regular point, then the two sequences above must converge to the same limit as $n\to\infty$. However, according to [9, Lemma 4.12], this common limit, whenever it exists, is a number $c>0$ independent of $\omega$. Therefore all LP-regular points of our Walters cocycle have the same Lyapunov exponents $c$ and c-c.